If the sum of two numbers is 2 and their difference is 1, what is their product.
Answer by John10(297)

If the sum of two numbers is 2 and their difference is 1, what is their product.
------------------------------------------------
Let x and y (larger number) be the numbers
x + y = 2
y - x = 1
Add two equations:
2y = 3
y = 3/2 = 1.5
If y = 1.5 then
x +1.5 = 2
x = 0.5
So x = 0.5 and y = 1.5 then their product is x*y = (0.5)(1.5) = 0.75
Hope it helps:)
John10